Things around you are emitting infrared radiation that includes the wavelength 1.03 x 10-5 m. What is the energy of the photons? ev
Answer #1

Solution

Let us go to the basics first.

We shall use following to find Energy E

E =h*f.....EQN.1


To find the frequency f, we shall use

f = C/λ, λ = 1.03 x10^ -5 meters
f = 3x 10^8/1.03 x 10^-5 = 2.912 x 10^13 Hertz

Now, h = Planck's constant = 4.136 x10^-15 electron volts per second

Thus, using EQN.1
E = (4.136 x 10^-15)(2.912 x 10^13) = 0.12 electron volts (eV) (Answer)

Thanks!!!